In​ randomized, double-blind clinical trials of a new​ vaccine,rats were randomly divided into two groups. Subjects in group 1 received the new vaccine while subjects in group 2 received a control vaccine. After the second​ dose,126 of 394 subjects in the experimental group​ (group 1) experienced drowsiness as a side effect. After the second​ dose, 170 of 620 of the subjects in the control group​ (group 2) experienced drowsiness as a side effect. Does the evidence suggest that a higher proportion of subjects in group 1 experienced drowsiness as a side effect than subjects in group 2 at the α=0.01 level of​ significance?

Find the test statistic?

Find the P-value?
Answer #1

Solution. Giren- Coup 2= 126 ni= 394 Group2 X2= 170 n2= 620 x=0.01 The null and alternative hypothesis is the Hoi p- f = 0 ts(0.3197 -0.2741) (0.291g (1-0.2919) (3 + 2) Zo= 1.5568 Zo=1.56) P-value- p-value = P(Z >Zo) =)-pczczo) =J-P(ZC).5568) =) - 0.